maximum value (n+l+m) for an unpaired electron in excited state of chlorine is

Grade:12

3 Answers

Aarti Gupta
10 years ago
Atomic no. of chlorine = 17
Ground state configuration = 1s2 2s2 2p6 3s2 3p5 ,in excited state the configuration become
3s1 3p6, thus,
For unpaired electron value of n = 3 ,l = 0 ,m = 0 so n+l+m = 3+0+0 = 3
Thanks & regards
Aarti gupta
askiitians faculty
simran kohli
10 years ago
but the answer is 7 not 3 @aarti
Kp99
7 years ago
The electron from P orbital of third shell will enter d orbital of the third shell so maximum value of n,l,m will be 2,3,3 resp. So n+L+m = 3+2+2=7